# Comparing zero-shot self-explanations with human rationales in multilingual text classification

### v1

@@ 4.4 Languages
4.5 Selected tokens
[…]
4.6 POS analysis
Figure 5: Top-6 POS tags of extracted rationales from SST monolingual for the human annotated rationales (left), self-explanations by llama3 (center) and post-hoc explanations based on llama3.
In order to gain a deeper insight into the differences between rationale selection, we are extracting part-of-speech (POS) tags of each selected tokens. We show results for the monolingual SST dataset in Figure 5. We see that for humans and self-explanations adjectives are most common, closely followed by nouns, verbs and with a bigger gap adverbs. For the post-hoc rationales, nouns were most often selected followed by verbs and adjectives. This means, the top-3 POS tags are the same across the different selection methods but differ in ranking between post-hoc explanations and the others. We see similar effects across models and languages, i.e., there is higher agreement in POS rankings between humans and self-explanations in comparison to post-hoc rationales. There are a few variations in ranking across languages, which has been seen in eye-tracking fixation times and attention scores before Brandl and Hollenstein (2022).
llama2 llama3 mistral mixtral
#tokens 0.11 0.04 0.34 0.16
json-syntax 0.86 0.02 0.05 0.37
4.7 Instruction following
4.8 Contrastive LRP
SST Refer to caption
RaFoLa Refer to caption
Comparing Cohen’s Kappa scores shown in Figure 6 suggests that contrastive post-hoc approaches do not generally result in higher plausibility than non-contrastive ones. Similarly, contrastive explanations overall exhibit a similar level of faithfulness as non-contrastive explanations (cf. Figure 4). Although contrastive explanations can be more faithful and plausible in certain cases, such as SST and mSST (English) for Llama3, there is no consistent difference between the two approaches. This aligns with previously reported high correlation between them Eberle et al. (2023) and the minor observed performance differences for generating the correct label based on contrastive or non-contrastive post-hoc approaches Krishna et al. (2023).
@@ 5 Discussion
In this paper, we investigated explanations generated by 4 different instruction-tuned LLMs for 2 text classification tasks in English but also in Italian and Danish. Those generated explanations, i.e., self-explanations, were limited to the input tokens of the respective text samples. We compared those self-explanations on the tasks of sentiment classification (SST) and forced labour classification (RaFoLa) with human annotations for the same samples and post-hoc feature attribution with layer-wise relevance propagation (LRP). Pairwise comparison between the three different types of explanations (humans, generated, post-hoc) shows that human annotations and generated explanations agree on a much higher level than post-hoc with any of them. We further saw that Llama3 has shown the highest level of agreement across both tasks and all languages and followed the instructions more closely in comparison to the other models. Our POS analysis confirmed this finding and showed higher agreement with the “type” of tokens selected by humans compared to post-hoc rationales.
@@
Besides established post-hoc attribution approaches, the ability of language models to provide self-generated explanations, i.e., self-explanations, has offered a direct and human-understandable communication between user and model. This not only enhances usability in particular for lay people but, as presented here, also results in explanations that are similarly faithful and align more closely with human rationales compared to binarized post-hoc attributions. Herein, the generation process behind self-explanations remains obfuscated and may suffer from counterfactuality, enabling the model to give untruthful explanations for correct predictions Ji et al. (2023). We further find that current language models require careful instructions to provide useful self-generated explanations.
While human plausibility is a desired property of predictions and explanations made by machine learning systems, LLMs are capable of identifying alternative solutions to task-solving that may not be intuitive to humans. Good explanations should thus also highlight the learned prediction strategy that is faithful to the one used by the model–even if it is not directly plausible Agarwal et al. (2024). In contrast to post-hoc attributions, which are directly derived from the prediction score (e.g., using explanatory gradients as for the LRP attributions considered here), the self-generation process is not explicitly tied to the prediction in a clear manner.
Attribution approaches herein consider the entire input to rank relevant features, including both the provided context and all instructions. Instruction templates are crucial for understanding and solving tasks, and we find that tokens related to task instructions often play a more significant role than the context itself. While this effectively identifies relevant task features from the perspective of the model, users are typically more interested in understanding which parts of the provided evidence are most relevant. Disentangling task instructions from contextual evidence is essential for aligning insights with user expectations and should inform the development of future interpretability methods.
Our study represents a first step toward understanding and building more intuitive model explanations by directly comparing human annotations with those generated by models. Evaluating free-text explanations for factuality, usability, and faithfulness is crucial for ensuring their practical and intuitive application, especially given the growing use of increasingly complex LLMs by lay people who may not fully understand their mechanisms
@@
We currently do not know why self-explanations align more closely with human annotations than post-hoc attributions. While training procedures such as reinforcement learning from human feedback Ziegler et al. (2019) may incentivize more human-like explanations Agarwal et al. (2024), limited access to models, procedures, and datasets, restricts detailed analysis.
The ability of the models to solve the sentiment classification task in unseen languages like Danish is remarkable. Not only is the model able to understand the Danish prompt, it also is able to return specific input tokens that are aligned with the human selection. Future work could investigate whether this zero-shot performance is specific to certain language families and tasks or if it extends to more complex instructions and datasets.
Conclusion As language models have advanced and become widely adopted, understanding the mechanisms behind their predictions has emerged as a critical challenge. Using human rationale annotations across three datasets, we have investigated the human plausibility and faithfulness of self-generated explanations to post-hoc feature attributions. We find that self-generated explanations agree more closely with human rationales than post-hoc feature attribution, while remaining faithful to the model prediction.
@@ Appendix B F1-Plausibility scores
Appendix C Most frequent rationale tokens

### v2

@@ 4.3 Faithfulness
Figure 5: Rationale token analysis of RaFoLa. (Left) Most frequent tokens extracted from the RaFoLa corpus, human annotations, and rationales extracted from model self-explanations and post-hoc explanations for articles #1 and #8. (Right) Ranking of named entities in articles class #1 extracted from model self-explanations and post-hoc explanations for Llama3. Resulting distributions are compared to the entity distribution found in human annotations.
5 Analyses
In this section, we provide a qualitative analysis of the differences in token selection across languages/risk indicators. We carry out part-of-speech/entity analyses and look into the most frequently selected tokens.
5.1 Sentiment Classification (SST)
POS We apply Part-of-Speech (POS) tagging to the selected tokens for SST/mSST. [^7] We use POS-tagging models from spaCy. We use human annotations as a baseline and analyse the difference in relative distribution over POS tags between the human baseline and self-explanations as well as LRP post-hoc rationales. Results for Mistral are shown in Figure 4, positive values signify relatively higher frequency than in human annotations.
The most important tokens include adjectives (ADJ), nouns (NOUN) and verbs (VERB) followed by adverbs (ADV). Both ADJ and ADV appear much less among selected tokens for both LRP as well as self-explanations in comparison to humans (negative values), with an exception for the English self-explanations (positive values). For Danish and Italian, ADJ/ADV appear even 10%/5% less, in post-hoc LRP than in the human annotations. On the other hand, pronouns (PRON) seem to be more relevant in LRP for English compared to human rationales. [^8] Please note, that we report percentage points here. We further see that nouns are used more across all languages in self-explanations in comparison to human annotations.
For the mSST dataset we see an increased usage of coordinating conjunction (CCONJ), i.e., words like and, but, for in the human annotations which then decreases across all languages for the models. This is most likely an effect of the difference in annotation guidelines for SST and mSST. [^9] Annotators in the SST study were asked to annotate individual words whereas guidelines for mSST required rather precise and connected phrases than individual tokens.
For Llama3 (Figure 12 in the Appendix), we see overall similar results with the main difference in the selection of adjectives in self-explanations. Across all languages we see an increase in the selection of adjectives for self-explanations in comparison to human rationales by about 5%, while the decrease for LRP is similar to the one in Mistral.
Selected Tokens We extract the top-8 most frequent tokens in the SST/mSST corpus. They are shown in Table 4 in the Appendix. Although the significance of those tokens is limited, we can see various trends that support the findings from our POS analysis in the previous section. Across models, we see that rationales in Mistral contain less adjectives and adverbs than for Llama2 and Llama3 in both generated and post-hoc rationales. We also overall see that across all models post-hoc rationales include less adjectives and adverbs than generated self-explanations where Llama2 relies the most on adjectives and adverbs.
5.2 Forced Labour detection (RaFoLa)
Selected tokens
To better understand the differences between model-based rationales and human annotations in detecting indicators of forced labor, we present in Figure 5 (left) the most frequent tokens for indicators #1 and #8, which show the lowest and highest agreements (see Section 4.2). We show other indicators in Table 3 in the Appendix.
When comparing the most frequent tokens as selected by human annotators and different models, we overall find that there is agreement in commonly used tokens, e.g. descriptive nouns covering the general topic of the dataset (workers, work, labour, etc.). Reflecting the shared use of natural language, human annotations and self-explanations show greater overlap in POS selection compared to post-hoc rationales (Figure 13 in the Appendix).
Comparing the indicators, article #8 (Physical and sexual violence) deviates more from the corpus’s most frequent tokens (first row) than #1. Keywords for #8, such as sexual and women, are more clearly identifiable, whereas keywords for #1 (Abuse of Vulnerability) may be harder to detect, possibly due to a less clear definition for an untrained model. This supports our finding that model probability scores are more affected by masking rationales in #8 than for #1 (Section 4.3).
Named Entities
By examining the distribution of named entities across rationale types, we next aim to uncover how annotations extracted from self-explanations and post-hoc explanations differ from humans rationales. Entities are extracted using spaCy and manually verified for correctness. In Figure 5 (right), we show the ranking of named entities across model-based rationales in Llama3 for article #1. Positive/negative scores indicate higher/lower occurrence of a specific entity type as compared to the entity distribution extracted from human annotations. We find that model-based rationales exhibit a stronger emphasis on organizations (ORG) than human rationales ($10-15\%$ more). Compared to models, human rationales instead prioritize broader societal categories, with increased usage of tokens referring to nationalities, religious or political groups (NORP, around $+5\%$), and geopolitical entities (GPE, $+3-8\%$). Post-hoc rationales consistently identify a greater number of person names (PER, $+4\%$) than both humans and self-explanations, as also shown in additional analyses in Appendix E.3 across models and indicators.
@@ 6 Discussion
In this paper, we evaluated self-explanations, i.e., explanations generated by instruction-tuned LLMs, based on their plausibility to humans and their faithfulness to models. We instructed 4 LLMs: Llama2, Llama3, Mistral and Mixtral for 2 text classification tasks in English but also in Italian and Danish. We constrained self-explanations to the input tokens of the respective text samples for which established evaluation methods can be easily applied. We analysed the sentiment classification (SST/mSST) and forced labour classification (RaFoLa) for which human annotations are available and included post-hoc feature attribution with layer-wise relevance propagation (LRP) for comparison.
Pairwise comparison between the 3 different types of explanations (humans, generated, post-hoc), i.e., plausibility, shows that human annotations and generated explanations agree much more strongly than post-hoc explanations with either. We further observe that Llama3 shows the highest level of agreement across both tasks and all languages. Our POS analysis further reveals a similar pattern of selecting rationales in SST/mSST between humans and self-explanations where both are mostly selecting adjectives, nouns and verbs. Across tasks, we see more pronounced differences for LRP, selecting less adjectives and adverbs for SST and more nouns in forced labour detection than humans. Fluctuations in human-model agreement for different POS tags, also across languages, have been observed before Brandl and Hollenstein (2022). The distribution of named entities in RaFoLa annotations reveals differences between models and humans, with models emphasizing organizations and person names, while humans annotate more geopolitical and religious entities, suggesting distinct approaches to collecting and evaluating evidence in forced labor detection.
Zero-shot performance in detecting forced labor varies across indicator types, with higher accuracies also resulting in higher agreements to human rationales for indicators #5 and #8. In comparison, the ability of models to solve the sentiment classification task in unseen languages like Danish is remarkable. Not only are the models able to understand the Danish prompt, they also return specific input tokens that are aligned with human rationales. We found that all models were able to make accurate predictions on both Danish and Italian, though it remains unclear whether this is due to data contamination during pre-training, e.g., training data for Llama2 has been reported to contain $0.11\%$ Italian Touvron et al. (2023).
@@
Faithfulness is a key desired property when evaluating the reliability of XAI methods. In our analyses, we found comparable faithfulness scores for self-explanations and post-hoc LRP that suggest that self-explanations could be a more accessible alternative to computation-intensive post-hoc explanations. Besides established post-hoc attribution approaches, the ability of language models to provide self-explanations has offered a direct and human-understandable communication between user and model. This not only enhances usability in particular for lay people but, as presented here, also results in explanations that are similarly faithful and align more closely with human rationales compared to binarized post-hoc attributions. Yet, the generation process behind self-explanations remains obfuscated and may suffer from counterfactuality, enabling the model to give untruthful explanations for correct predictions Ji et al. (2023). In particular, predictions and explanations made by LLMs can identify alternative solutions to task-solving that may not be intuitive to humans. Therefore, good explanations should highlight the learned prediction strategy, ensuring it is faithful to the model’s approach, even if it is not directly plausible Agarwal et al. (2024).
@@
In summary, we find higher plausibility for self-explanations compared to post-hoc rationales while maintaining faithfulness. However, for more challenging tasks, we still see room to improve the plausibility of self-explanations. We further see that LLMs require careful instructions to generate useful self-explanations. We currently do not know why self-explanations align more closely with human annotations than post-hoc attributions. While training procedures such as reinforcement learning from human feedback Ziegler et al. (2019) may incentivize more human-like explanations Agarwal et al. (2024), limited access to models, procedures and data, restricts detailed analysis.
Our study represents a first step toward understanding and building more intuitive model explanations by directly comparing human annotations with those generated by models. Evaluating free-text explanations for factuality, usability, and faithfulness is crucial for ensuring their practical and intuitive application, especially given the growing use of increasingly complex LLMs by lay people who may not fully understand their mechanisms.
@@ Appendix B F1-Plausibility scores
Table 2: Ratios of identified rationale tokens by humans and those of self-generated rationales across models for SST and RaFoLa.
ratio ratio model
human llama llama3 mistral mixtral
SST 0.29 0.22 0.25 0.32 0.14
mSST (EN) 0.33 0.23 0.27 0.32 0.15
mSST (DK) 0.38 0.18 0.29 0.40 0.21
mSST (IT) 0.37 0.26 0.31 0.39 0.25
RaFoLA #1 0.05 0.05 0.23 0.12 0.16
RaFoLA #2 0.06 0.04 0.21 0.11 0.12
RaFoLA #5 0.04 0.02 0.08 0.06 0.05
RaFoLA #8 0.07 0.04 0.12 0.08 0.10
Appendix C Rationale ratios
Table 2 presents ratios of identified rationale tokens for SST, multilingual SST (mSST), and RaFoLa datasets in humans and models. The ratio of identified rationale tokens by humans tends to be higher for the shorter movie reviews in SST and mSST, and lower for the longer paragraphs in RaFoLa, suggesting that the human annotators have more sparsely identified relevant phrases when presented with longer texts. This can be attributed to the differences in samples across the datasets. While movie reviews in SST tend to be short and were written with the purpose of expressing a sentiment, the articles in RaFoLa were more descriptive of a specific situation or incident and might or might not include the violation of one (or more) of the 11 risk indicators.
Figure 12: Distribution of POS-tags in comparison to top-6 POS in human annotations for SST. Absolute POS tags have been normalized (so they sum to 1) prior to computing the difference to human baseline. Ranking is based on prevalence in human rationales. Difference is shown for self-explanations (left) and post-hoc LRP (right) across all data splits and languages for Llama3.
Figure 13: Distribution of POS-tags in comparison to top-6 POS in human annotations for RaFoLa. Absolute POS tags have been normalized (so they sum to 1) prior to computing the difference to human baseline. Ranking is based on prevalence in human rationales. Difference is shown for self-explanations and post-hoc LRP across all data splits and languages for Mistral, (left) and Llama3 (right).
Appendix D POS analysis
Figure 12 shows difference in relative POS distribution between model rationales and humans for Llama3 on SST/mSST. Positive values signify higher frequency in comparison to human rationales. Results for this and a similar figure for Mistral (Figure 4) are discussed in Section 5.1. Figure 13 show the same POS analysis for RaFoLa.
Appendix E Most frequent rationale tokens
@@
llama2 llama3 mistral mixtral
#tokens 0.11 0.04 0.34 0.16
json-syntax 0.86 0.02 0.05 0.37
E.1 Instruction following
E.2 Contrastive post-hoc explanations
SST Refer to caption
RaFoLa Refer to caption
SST
mSST      English
Danish
Italian
RaFoLa        #1
#2
#5
#8
Figure 15: Faithfulness evaluation for SST and mSST (top row) and RaFoLa (bottom row). The probability after masking the tokens extracted from human rationale annotations, generated model rationales (self-explanations) and post-hoc model attributions is compared across models. Lower probability indicates more faithful identification of rationales.
Comparing Cohen’s Kappa scores shown in Figure 14 suggests that contrastive post-hoc approaches do not generally result in higher plausibility than non-contrastive ones. Similarly, contrastive explanations overall exhibit a similar level of faithfulness as non-contrastive explanations (cf. Figure 15). Although contrastive explanations can be more faithful and plausible in certain cases, such as SST and mSST (English) for Llama3, there is no consistent difference between them. This aligns with previously reported high correlation between them Eberle et al. (2023) and the minor observed performance differences for generating the correct label based on contrastive or non-contrastive approaches Krishna et al. (2023).
E.3 Entity analysis
Figure 16: Entity analysis of Rafola for #1 (top) and #8 (bottom) for llama3 (left) and mistral (right), showing top-8 entities across rationale types (human, model, post-hoc).